# Time Transfer: On Optimal Learning Rate and Batch Size In The Infinite Data Limit

### v1

@@ Abstract
One of the main challenges in optimal scaling of large language models (LLMs) is the prohibitive cost of hyperparameter tuning, particularly learning rate $\eta$ and batch size $B$. While techniques like $\mu$(Yang et al., 2022) provide scaling rules for optimal $\eta$ transfer in the infinite model size limit, the optimal scaling behavior in the infinite data size limit ($T\to\infty$) remains unknown. We fill in this gap by observing for the first time an interplay of three optimal $\eta$ scaling regimes: $\eta\propto\sqrt{T}$, $\eta\propto 1$, and $\eta\propto 1/\sqrt{T}$ with transitions controlled by $B$ and its relation to the time-evolving critical batch size $B_{\mathrm{crit}}\propto T$. Furthermore, we show that the optimal batch size is positively correlated with $B_{\mathrm{crit}}$: keeping it fixed becomes suboptimal over time even if learning rate is scaled optimally. Surprisingly, our results demonstrate that the observed optimal $\eta$ and $B$ dynamics are preserved with $\mu$P model scaling, challenging the conventional view of $B_{\mathrm{crit}}$ dependence solely on loss value. Complementing optimality, we examine the sensitivity of loss to changes in learning rate, where we find the sensitivity to decrease with $T\to\infty$ and to remain constant with $\mu$P model scaling. We hope our results make the first step towards a unified picture of the joint optimal data and model scaling.
@@ 1 Introduction
Large Language Models (LLMs) have increasingly become a prominent area of study in the field of Natural Language Processing (NLP) and beyond. They have demonstrated significant improvement in performance across a wide range of tasks, such as language understanding, text generation, translation, and summarization, showing results comparable or outperforming those of an average domain expert (Dubey et al., 2024; OpenAI et al., 2024; Team et al., 2023). The primary advantage of LLMs is their ability to scale well with increased computational resources, which results in predictive improved performance (Kaplan et al., 2020; Hoffmann et al., 2022).
@@
  • Optimal learning rate scaling: depending on batch size $B$ and point in time during the training, we observe three regimes of the optimal learning rate $\eta^{*}$ evolution in the limit of increasing pretraining token budget $T\to\infty$ (Fig. 1(a)):
    • $\eta^{*}\propto 1/\sqrt{T}$ for $B<B_{\mathrm{crit}}$,
    • $\eta^{*}\propto 1$ for $B\sim B_{\mathrm{crit}}$,
    • $\eta^{*}\propto\sqrt{T}$ for $B>B_{\mathrm{crit}}$,
    where transitions between them are controlled by $B$ relative to the critical batch size $B_{\mathrm{crit}}$ (see Sec. 2.1 for definition and Appendix A.2 for general discussion). The region boundaries are dynamically evolving in time following the critical batch size evolution $B_{\mathrm{crit}}(T)$. Furthermore, we find these dynamics to be largely preserved within $\mu$P (Appendix A.9).
  • Optimal batch size scaling: assuming $\eta$ is optimal for a given data horizon $T$, we observe an approximately $B^{*}\propto\sqrt{T}$ drift of the optimal batch size $B^{*}$ (Fig. 2(a)). The drift is correlated with the drift of the $B_{\mathrm{crit}}$ region, with $B^{*}$ falling within the region. Importantly, we show that naïve application of optimal $\eta$ scaling rules in the $T\to\infty$ limit with $B$ being indefinitely fixed becomes suboptimal over time: a joint $(\eta,B)$ scaling is required.
  • Critical batch size: $B_{\mathrm{crit}}$ evolves in time with approximately $B_{\mathrm{crit}}\propto T$ linear dependency on the training token budget (Fig. 1(b)). This dynamics also affects optimal $\eta$ scaling regimes, driving the time transition between them. Surprisingly, we show evidence that $B_{\mathrm{crit}}$ is not exclusively defined by the value of the loss function (Eq. 7) as suggested by McCandlish et al. (2018): models within $\mu$P share the same $B_{\mathrm{crit}}$ region while having different performance in terms of loss.
@@ 2.1 Terminology
Critical batch size $\boldsymbol{(B_{\mathrm{crit}})}$: similarly to the SGD study of Shallue et al. (2019), we define $B_{\mathrm{crit}}$ as the region where the $\eta\propto\sqrt{B}$ scaling rule for a fixed token budget breaks and, as we observe, directly switches to a $\eta\propto 1/\sqrt{B}$ scaling rule. This corresponds to the peak of the bell-shaped curve (Fig. 1(b)), which was shown by Li et al. (2024) to equal the $B_{\mathrm{crit}}$ definition of McCandlish et al. (2018) (see Appendix A.2 for extended discussion).
@@ 2.2 Model configuration and datasets
For all our experiments we use a default MPT model architecture (MosaicML, 2023) as implemented in the llm-foundry codebase (MosaicML, 2024), with all the models sharing the same training configuration (Appendix A.3). We use the Decoupled AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e^{-8}$, weight decay $\lambda=0$ and gradient clipping by the $L_{2}$ norm value of 1.
@@ 3.1 Learning rate optimum drifts in time, with batch size interpolating between different scaling rules
Figure 1: (a) Evolution of the optimal learning rate with an increase of the pretraining token budget $\eta^{*}_{\mathrm{norm}}(T)$ , normalized to $\eta^{*}|_{T=2^{30}}$, for a set of batch sizes (in tokens). Each point is obtained by averaging optimal learning rate values across $\mu$P model family, as described in Sec. 3.1. Dashed lines correspond to square-root $\eta^{*}\propto\sqrt{T^{(-1)}}$ scaling rules. We observe an $\eta^{*}\propto 1/\sqrt{T}$ regime for batch sizes $B\in\{2^{16},2^{18}\}$ saturating after $T=2^{35}$, an intermediate regime $\eta^{*}\propto 1$ for $B\in\{2^{20},2^{22}\}$, and an $\eta^{*}\propto\sqrt{T}$ regime for $B\in\{2^{24},2^{26}\}$. (b) Transposition of Fig. 1(a): optimal learning rate $\eta^{*}$ per batch size, against a range of pretraining token budgets. Each point is $\mu$P-averaged as in (a), with color bands visualizing the corresponding standard deviation. Arrows indicate the peak batch size value for each of the fixed budget curves, referred to as critical batch size $B_{\mathrm{crit}}$ (Sec. 2.1). We observe an approximately linear evolution of $B_{\mathrm{crit}}$ in the limit of increased token budget, not present in the $\mu$P width limit (Appendix  A.9).
[…]
Fig. 1(a) illustrates how batch size serves as a parameter interpolating between various learning rate scaling regimes. For the smallest probed batch sizes $B=\{2^{16},2^{18}\}$ we observe an initially perfect inverse square-root scaling of the optimal learning rate $\eta^{*}$ with increase of the data horizon $T$. The scaling, however, breaks down at some point and optimal learning rate plateaus. We speculate that this may be related to gradient moments accumulation and the fact that we apply no scaling rules for the $\beta_{1,2}$ AdamW parameters, in order to reduce analysis complexity. Such rules were derived by Hilton et al. (2022); Malladi et al. (2023), although only in the fixed data budget setup. Understanding the ways of extending this scaling regime further in time provides an interesting direction for future studies.
The $\eta^{*}$ drift in time is not per se a new phenomenon: a linear rule for SGD (Smith et al., 2020) and a square-root rule for Adam (Shen et al., 2024) are already known. We complement this picture with observing, for the first time, richer dynamics: namely, a continuous transition from inverse square-root to square-root scaling regimes. As the batch size increases, one first observes (inverse) constant-to-moderate scaling for $B=2^{20}(2^{22})$ which we also refer to as a metastable regime: to a certain degree, optimal learning rate transfers in time before breaking into one of the square-root regimes. Then, the metastable regime changes to square-root scaling for larger batch sizes $B=\{2^{24},~{}2^{26}\}$. The scaling does not appear immediately after the beginning of the training but rather after some time. We link this with the effect of “lost” epochs (Smith et al., 2018), which states that larger batch sizes require a larger number of iterations for the gradient accumulation to reach steady convergence dynamics. In our case, the transition to the square-root scaling happens only after iterating through $2^{34}\text{--}2^{35}\approx 17\text{--}34\mathrm{B}$ tokens and takes up to $137\mathrm{B}$ tokens to establish.
@@ 3.2 Optimally-tuned batch size increases in time
We observe an approximate square-root scaling of the optimal batch size with increase of the pretraining token budget from $B^{*}|_{T=2^{30}}=2^{18}$ to $B^{*}|_{T=2^{35}}=2^{20}$ (Fig. 2(a)). Emergence of suboptimality is more pronounced when transposing the token budget and batch size axes (Fig. 2(b)), where the smallest $B=2^{16}$ batch size curve, with each point having learning rate scaled with the inverse scaling rule $\eta^{*}\propto 1/\sqrt{T}$, is being taken over in the $T\rightarrow\infty$ limit by the curves corresponding to larger batch sizes.
@@
This result illustrates that, while naïve “pairwise” scaling rules for optimal learning rate, e.g. $\eta^{*}\propto 1/\sqrt{T}$, are convenient for predicting optimal values at scale, they do not necessarily result in the best model performance: taking batch size dynamics into account is required. In other words, the invariant induced solely by the $\eta^{*}\propto 1/\sqrt{T}$ scaling rule is not sufficient for the model performance to be optimal. We believe, similarly to Smith & Le (2018), that some broader notion of noise scale should serve as a more fundamental invariant to optimize for in the joint data and model size limit. We discuss this idea in more detail in Sec. 4.
3.3 Critical batch size region evolves in time, but is unchanged within $\mu$P
In Fig. 1(b), we reinterpret Fig. 1(a) by transposing batch size and token budget axes and plotting the optimal learning rate and batch size scaling jointly per data horizon. This better emphasizes dynamics in the fixed token budget regime, rather than the trajectory of optimal learning rate evolution in time. As in Sec. 3.1, we perform an average across $\mu$P models sharing the same $d_{\mathrm{model}}^{\mathrm{base}}=1024$ and also plot the corresponding standard deviation. We include a similar plot for the other base model with $d_{\mathrm{model}}^{\mathrm{base}}=256$ in Appendix A.8 and individual plots for each of the $(d_{\mathrm{model}},~{}d_{\mathrm{model}}^{\mathrm{base}})$ configurations in Appendix A.9.
We observe that for a given time horizon, the $(\eta^{*},B)$ curve has a bell-like shape, as predicted by Li et al. (2024). The left-hand side of the peak represents a previously known $\eta\propto\sqrt{B}$ scaling rule (Malladi et al., 2023; Shen et al., 2024). However, with our experiments, we uncover a previously unseen right-hand side of the curve, also referred to as “surge” by Li et al. (2024), where the optimal learning rate for a fixed token budget scales inversely proportionally to the batch size scaling via the $\eta^{*}\propto 1/\sqrt{B}$ rule. Although statistical uncertainties are high for the experiment with $d_{\mathrm{model}}^{\mathrm{base}}=1024$, the square-root rules are more pronounced for the experiment with $d_{\mathrm{model}}^{\mathrm{base}}=256$ (Appendix A.8).
Furthermore, the peak position of the fixed token budget, which we refer to as the critical batch size $B_{\mathrm{crit}}$ (Sec. 2.1), is evolving in time via an approximately $B_{\mathrm{crit}}\simeq T$ scaling rule, showing the same trend as in McCandlish et al. (2018). Moreover, we note that the optimal learning rate corresponding to $B_{\mathrm{crit}}$ decreased by a factor of two with a horizon scaling from $2^{30}$ to $2^{37}$ tokens. This might be a statistical fluctuation since there is no significant decrease for the experiment with $d_{\mathrm{model}}^{\mathrm{base}}=256$ (Appendix A.8).
Lastly, there is a difference of $B_{\mathrm{crit}}$ evolution between the $T$ and $\mu$P infinite width limits. Specifically, for a fixed token budget, we observe no significant change of the curves’ shapes and peak positions across $d_{\mathrm{model}}$ values within the same $\mu$P trajectory, and also with the change of the base model (Appendix A.9). At the same time, there is a noticeable drift of $B_{\mathrm{crit}}$ in the $T\to\infty$ limit with the model being fixed. As both limits are accompanied with a comparable change of the model performance [^2] Back-of-the-envelope calculation from Fig. 2(a) and Appendix A.10: for $d_{\mathrm{model}}^{\mathrm{base}}=1024$, $B=2^{20}$, there is a loss change $\mathcal{L}_{\mathrm{val}}=3.4\to 2.8$ with a token budget increase $2^{31}\to 2^{37}$, resulting in $B_{\mathrm{crit}}$ drifting by $2^{4}$. For the same $(d_{\mathrm{model}}^{\mathrm{base}},B)$ configuration, there is no significant $B_{\mathrm{crit}}$ drift with a change of width by $2^{2}$ within $\mu$P, but the corresponding loss change is $\mathcal{L}_{\mathrm{val}}=3.5\to 2.9$., this observation brings evidence that dependence of the critical batch size exclusively on the loss value suggested by Kaplan et al. (2020) (Eq. 7) is not entirely complete. Or, contrary to experimental results in Li et al. (2024), the two definitions of the critical batch size region (Appendix A.2) are not the same and should be disentangled.

### v2

@@ Abstract
One of the main challenges in optimal scaling of large language models (LLMs) is the prohibitive cost of hyperparameter tuning, particularly learning rate $\eta$ and batch size $B$. While techniques like $\mu$(Yang et al., 2022) provide scaling rules for optimal $\eta$ transfer in the infinite model size limit, the optimal scaling behavior in the infinite data size limit remains unknown. We fill in this gap by observing for the first time an intricate dependence of optimal $\eta$ scaling on the pretraining token budget $T$, $B$ and its relation to the critical batch size $B_{\mathrm{crit}}$, which we measure to evolve as $B_{\mathrm{crit}}\propto T$. Furthermore, we show that the optimal batch size is positively correlated with $B_{\mathrm{crit}}$: keeping it fixed becomes suboptimal over time even if learning rate is scaled optimally. Surprisingly, our results demonstrate that the observed optimal $\eta$ and $B$ dynamics are preserved with $\mu$P model scaling, challenging the conventional view of $B_{\mathrm{crit}}$ dependence solely on loss value. Complementing optimality, we examine the sensitivity of loss to changes in learning rate, where we find the sensitivity to decrease with increase of $T$ and to remain constant with $\mu$P model scaling. We hope our results make the first step towards a unified picture of the joint optimal data and model scaling.
@@ 1 Introduction
Large Language Models (LLMs) have increasingly become a prominent area of study in the field of Natural Language Processing (NLP) and beyond. They have demonstrated significant improvement in performance across a wide range of tasks, such as language understanding, text generation, translation, and summarization, showing results comparable or outperforming those of an average domain expert (Dubey et al., 2024; Achiam et al., 2023; Team et al., 2023). The primary advantage of LLMs is their ability to scale well with increased computational resources, which results in predictive improved performance (Kaplan et al., 2020; Hoffmann et al., 2022).
@@
  • Optimal learning rate scaling: by incorporating a dependence on the pretraining token budget into the theoretical model of optimal learning rate $\eta^{*}$ scaling Li et al. (2024) via Eq. 1 and performing a fit to experimentally observed data (Fig. 1), we establish a dependence of the $\eta^{*}$ evolution with $T$ on the batch size $B$ and its relation to the critical batch size $B_{\mathrm{crit}}$ (see definition in Sec. 2.1). From interpreting the model fit results, we obtain scaling behaviors ranging from $\eta^{*}\propto\sqrt{T}$ to $\eta^{*}\propto 1/\sqrt{T}$ depending on $B$, $B_{\mathrm{crit}}$ and $T$, which we find compatible with experimental observations. Furthermore, we find these dynamics to be largely preserved within $\mu$P (Appendix A.11).
  • Optimal batch size scaling: assuming $\eta$ is optimal for a given data horizon $T$, we observe a gradual increase of the optimal batch size $B^{*}$ with an increase of the token budget (Fig. 3(a)). The drift is correlated with the evolution of the critical batch size $B_{\mathrm{crit}}$ (Fig. 2, left), with $B^{*}(T)<B_{\mathrm{crit}}(T)$ in our measured range of $T$. Importantly, we show that naïve application of optimal $\eta$ scaling rules in the $T\to\infty$ limit with $B$ being indefinitely fixed becomes suboptimal over time: a joint $(\eta,B)$ scaling is required.
  • Critical batch size: we experimentally find $B_{\mathrm{crit}}$ (see definition in Sec. 2.1) to evolve in time with $B_{\mathrm{crit}}\propto T^{\alpha_{B}}$ and $\alpha_{B}=1.0\pm 0.2$ (Fig. 2, left). This dynamic affects optimal $\eta$ scaling via Eq. 1 and drives the transition between various scaling behaviors (Sec. 3.2). Surprisingly, we show evidence that $B_{\mathrm{crit}}$ is not exclusively defined by the value of the loss function (Eq. 8) as suggested by McCandlish et al. (2018): models within $\mu$P share the same $B_{\mathrm{crit}}$ region while having different performance in terms of loss.
@@ 2.1 Terminology
Critical batch size $\boldsymbol{(B_{\mathrm{crit}})}$: following Li et al. (2024), we define $B_{\mathrm{crit}}$ as the corresponding parameter in Eq. 1, also describing the peak position of the bell-shaped curve (Fig. 1(a)), which was shown by the authors to equal the $B_{\mathrm{crit}}$ definition of McCandlish et al. (2018). To better clarify the nomenclature appearing in the literature, we provide an extended discussion in Appendix A.2.
@@ 2.2 Model configuration and datasets
For all our experiments we use a default MPT model architecture (MosaicML, 2023) as implemented in the llm-foundry codebase (MosaicML, 2024), with all the models sharing the same training configuration (Appendix A.3). We use the Decoupled AdamW optimizer (Loshchilov, 2017) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$, weight decay $\lambda=0$ and gradient clipping by the $L_{2}$ norm value of 1.
@@ 3 Results
3.1 Critical batch size evolves in time, but is unchanged within $\mu$P
Figure 1: (a): Optimal learning rate $\eta^{*}$ per batch size against a set of pretraining token budgets (see Appendix A.9 for a full set). Each point is obtained by averaging experimental observations of optimal learning rate values across $\mu$P model family and random seeds, as described in Sec. 3.1, with color bands visualizing the corresponding standard deviation. Solid lines represent the fitted theoretical model of Li et al. (2024) (Eq. 1) as described in Sec. 3.1, dashed lines only connect the data points for visualization purposes. We observe an approximately linear growth of $B_{\mathrm{crit}}$ (see also a dedicated Fig. 2), defined as the peak position of the fitted curve, in the limit of increased token budget.
(b) Transposition of Fig. 1(a): evolution of the optimal learning rate with an increase of the pretraining token budget $\eta^{*}(T)$ for a representative set of batch sizes, in tokens. We observe the fitted model to describe the scaling behavior of low ($B=2^{18}$) and high ($B=2^{26}$) batch sizes, as well as intermediate batch sizes in the high token budget regime. For $B=2^{18}$, the model reduces to $\eta^{*}\propto 1/\sqrt{T}$ as discussed in Sec. 3.2, matching the observations.
First, we begin with setting $d_{\mathrm{model}}^{\mathrm{base}}=1024$ and scanning learning rate across different batch sizes and $d_{\mathrm{model}}$. We present results for the $\eta^{*}$ optimum dependence on the batch size $B$ per data horizon $T$ in Fig. 1(a), with individual $\mathcal{L}_{\mathrm{val}}(\eta)$ profile scans in Appendix A.8 and the full set of horizons in Appendix A.9. In order to reduce statistical uncertainties, we average results across three $\mu$P models [^1] We believe this averaging approach is justified since all the three models share the same optimization trajectory in terms of the number of steps, batch size and data horizon length, therefore are theoretically guaranteed by $\mu$Transfer to share the same optimal learning rate. From the experimental side, we also observe no significant differences across the three models (Appendix A.11). with $d_{\mathrm{model}}\in\{256,~{}512,~{}1024\}$ for tokens budgets $T\leq 2^{35}$ and additionally across four more random seeds for large batch size values $B\in\{2^{20},~{}2^{22},~{}2^{24},~{}2^{26}\}$ for the model with $d_{\mathrm{model}}=256$ to reduce statistical fluctuations in the low token budget region. We include a similar plot for the other base model with $d_{\mathrm{model}}^{\mathrm{base}}=256$ in Appendix A.10 and individual plots for each of the $(d_{\mathrm{model}},~{}d_{\mathrm{model}}^{\mathrm{base}})$ configurations in Appendix A.11.
We observe that for a given time horizon, the $(\eta^{*},B)$ curve has a bell-like shape, as predicted by Li et al. (2024). The left-hand side of the peak represents a known $\eta\propto\sqrt{B}$ scaling rule (Malladi et al., 2022; Shen et al., 2024). However, with our experiments, we uncover a previously unseen right-hand side of the curve, also referred to as “surge” by Li et al. (2024), where the optimal learning rate for a fixed token budget scales inversely proportionally to the batch size scaling via the $\eta^{*}\propto 1/\sqrt{B}$ rule.
We analyze the observed data points within the theoretical framework of Li et al. (2024). For each of the token budgets we fit the data with the following functional form:
$\eta^{*}(T,B)=\dfrac{\eta_{\mathrm{crit}}(T)}{\sqrt{\frac{B}{B_{\mathrm{crit}}(T)}}+\sqrt{\frac{B_{\mathrm{crit}}(T)}{B}}},$ (1)
where $B_{\mathrm{crit}}$ and $\eta_{\mathrm{crit}}$ are parameters of the fit. The former corresponds to the peak position of the bell-like curve and was shown in McCandlish et al. (2018); Li et al. (2024) to approximate the critical batch size defined as the balance point between optimal number of training steps and data efficiency. The latter can be interpreted as the optimal learning rate when training in the regime with the batch size tracking the critical one, i.e. $B(T)=B_{\mathrm{crit}}(T)$.
After performing the fit for each token budget, we analyze the $\eta_{\mathrm{crit}}$ and $B_{\mathrm{crit}}$ evolution in time. We fit a power law of the form $p_{\mathrm{crit}}=a_{p}T^{\alpha_{p}}+b_{p},~{}p\in\{\eta,B\}$ to each of the data sets following the procedure described in Appendix A.5 and present the results in Fig. 2 (solid lines). For the time dependence of critical parameters, we obtain the following power exponents:
$B_{\mathrm{crit}}$ $\propto T^{\alpha_{B}},\hskip 14.22636pt\alpha_{B}=1.0\pm 0.2,$ (2)
$\eta_{\mathrm{crit}}$ $\propto T^{\alpha_{\eta}},\hskip 15.649pt\alpha_{\eta}=-1.3\pm 0.4,$
where we find that in both cases the hypothesis of the power exponent +1(1)11+1~{}(-1)+ 1 ( - 1 ) for $\alpha_{B}~{}(\alpha_{\eta})$ is compatible with experimental observations within uncertainties. Detailed results for the $a,b$ coefficients are provided in Appendix A.5.
Figure 2: Critical batch size $B_{\mathrm{crit}}$ (left) and critical learning rate $\eta_{\mathrm{crit}}$ (right), as extracted from the fit with the power law $p_{\mathrm{crit}}=a_{p}T^{\alpha_{p}}+b_{p}$, where $p\in\{\eta,B\}$, following the procedure from Appendix A.5, as a function of token budget. Solid line represents the fit result. Dashed line corresponds to the fit with the power exponent fixed to $\alpha_{B}=0.5$ (left) and $\alpha_{\eta}=-0.5$ (right). This model fit is visualized only to illustrate the model variation with the exponent change and its parameters are not used in the main analysis.
Lastly, there is a difference of $B_{\mathrm{crit}}$ evolution between the $T$ and $\mu$P infinite width limits. Specifically, for a fixed token budget, we observe no significant change of the curves’ shapes and peak positions across $d_{\mathrm{model}}$ values within the same $\mu$P trajectory, and also with the change of the base model (Appendix A.11). At the same time, there is a noticeable drift of $B_{\mathrm{crit}}$ in the $T\to\infty$ limit with the model being fixed. As both limits are accompanied with a comparable change of the model performance [^2] Back-of-the-envelope calculation from Fig. 3(a) and Appendix A.12: for $d_{\mathrm{model}}^{\mathrm{base}}=1024$, $B=2^{20}$, there is a loss change $\mathcal{L}_{\mathrm{val}}=3.4\to 2.8$ with a token budget increase $2^{31}\to 2^{37}$, resulting in $B_{\mathrm{crit}}$ drifting by $2^{4}$. For the same $(d_{\mathrm{model}}^{\mathrm{base}},B)$ configuration, there is no significant $B_{\mathrm{crit}}$ drift with a change of width by $2^{2}$ within $\mu$P, but the corresponding loss change is $\mathcal{L}_{\mathrm{val}}=3.5\to 2.9$., this observation brings evidence that dependence of the critical batch size exclusively on the loss value suggested by Kaplan et al. (2020) (Eq. 8) is not entirely complete. Or, contrary to results in Li et al. (2024), the two definitions of the critical batch size region (Appendix A.2) are not the same and should be disentangled.
@@ 3.2 Learning rate optimum drifts in time, with batch size interpolating between different scaling rules
In Fig. 1(b), we reinterpret Fig. 1(a) by transposing the batch size and token budget axes and by plotting the evolution of the optimal learning rate $\eta^{*}$ in time $T$ for a representative set of batch size values, with the full set of batch size values in Appendix A.9. Overlayed, we also plot the model fitted with Eq. 1 (solid lines).
From the data points alone we observe an intricate drift of the optimal learning rate in time as governed by the batch size value. In a simplified way, for small $B$ values ($2^{18}$ and $2^{20}$ in Fig. 1(b)), we observe a decrease of $\eta^{*}$ by $2^{2}$ with an increase of the token budget by $2^{7}$, while for the larger $B$ values ($2^{24}$ and $2^{26}$), it is oppositely an increase by up to $2^{2}$.
We also find that the fitted model [^3] We note that the model is not fitted to the data points in the representation of Fig. 1(b). Instead, as described in Sec. 3.1, the fit is performed in two consecutive steps: by first capturing the $(\eta^{*},B)$ behavior per token budget (illustrated in Fig. 1(a)), followed by fitting the time dynamics (illustrated in Fig. 2). describes the data points for the smallest ($2^{16}$ and $2^{18}$) and largest ($2^{26}$) probed batch sizes well. For the intermediate batch size values, it captures the behavior in the large token budget regime and the general curvature patterns ($B=2^{24}$), but sometimes lacks the correct amplitude. We note, however, large uncertainties on the fitted model parameters (Appendix A.5): additional data points with improved resolution in both $\eta$ and $B$ would better constrain the model fit and therefore constitute an important next step in future work.
It is instructive to consider several limiting scaling scenarios of Eq. 1. First, when $B\ll B_{\mathrm{crit}}$, one obtains $\eta^{*}(T,B)\propto\eta_{\mathrm{crit}}(T)/\sqrt{B_{\mathrm{crit(T)}}}$, which we do not observe, since $\min(B_{\mathrm{crit}})=2^{18}$ in our experiments. Second, when $B\gg B_{\mathrm{crit}}$, one obtains $\eta^{*}(T,B)\propto\eta_{\mathrm{crit}}(T)\cdot\sqrt{B_{\mathrm{crit(T)}}}$. We observe this regime for high batch size values ($B=2^{24}$ and $B=2^{26}$) in the high token budget regime ($T>2^{34}$ tokens). Since in this region, we see $\eta_{\mathrm{crit}}(T)\sim 1$ (Fig. 2, right) and $B_{\mathrm{crit}}\propto T^{\alpha_{B}}\approx T$, we obtain $\eta^{*}\propto\sqrt{T}$ (as can be seen in Fig. 1(b)). Lastly, a special case is when $B(T)=B_{\mathrm{crit}}(T)$, i.e. batch size is tracking the critical one. We observe this regime for the smallest batch sizes ($B=2^{16}$ and $B=2^{18}$) in the low token budget regime ($T<2^{33}$ tokens). In that case, the optimal learning rate $\eta^{*}\propto\eta_{\mathrm{crit}}\propto T^{\alpha_{\eta}}=T^{-0.5}$ (see Fig. 2, where the power exponent $\alpha_{\eta}=-0.5$ describes data in the low token budget region better).
@@ 3.3 Optimally-tuned batch size increases in time
We observe an increase of the optimal batch size with increase of the pretraining token budget from $B^{*}|_{T=2^{30}}=2^{18}$ to $B^{*}|_{T=2^{35}}=2^{20}$ (Fig. 3(a)). Emergence of suboptimality is more pronounced when transposing the token budget and batch size axes (Fig. 3(b)), where the smallest $B=2^{16}$ batch size curve, with each point having learning rate scaled approximately with the inverse scaling rule $\eta^{*}\propto 1/\sqrt{T}$, is being taken over in the $T\rightarrow\infty$ limit by the curves corresponding to larger batch sizes. Furthermore, comparing the optimal batch size values with the critical batch size evolution (Sec. 3.1), we obtain $B^{*}(T)<B_{\mathrm{crit}}$ for the range of our measurements $T=[2^{30},~{}2^{37}]$.
@@
This result illustrates that, while naïve “pairwise” scaling rules for optimal learning rate, e.g. $\eta^{*}\propto 1/\sqrt{T}$, are convenient for predicting optimal values at scale, they do not necessarily result in the best model performance: taking batch size dynamics into account is required. In other words, the invariant induced solely by, for example, the $\eta^{*}\propto 1/\sqrt{T}$ scaling rule is not sufficient for the model performance to be optimal. We believe, similarly to Smith & Le (2017), that some broader notion of noise scale should serve as a more fundamental invariant to optimize for in the joint data and model size limit. We discuss this idea in more detail in Sec. 4.